If AM and HM of the roots of a quadratic equation are 3/2 and 4/3 respetively, then that equation is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Abhishek
 
Let roots are A and B
 
Hence
 
(A+B)/2 = 3/2
 
A+ B = 3
 
Now
 
2AB/A+ B = 4/3
 
2 AB/3 = 4/3
AB = 2
 
 
Hence
Equation 
 
x² - 3x + 2 = 0
 
 
Regards
Arun (askIITians forum expert)